# Regularized KL-Divergence for Well-Defined Function-Space Variational Inference in Bayesian Neural Networks

### v2

@@ A.3.8 Hardware
A.4 Additional experimental results
@@
A.4.1 Qualitative uncertainty evaluation
@@
We further find that our method (GFSVI) provides strong regularization when the data generative process is noisy (see figure 4) and is more robust than FVI to situations where ones computational budget constrains the number of measurement points $M$ to be small (figure 8). In contrast to FVI, GFSVI accurately approximates the exact GP posterior under rough (Matern-1/2) GP priors effectively incorporating prior knowledge defined by the GP prior to the inference process (see figure 5). Likewise, GFSVI adapts to the variability of the functions specified by the kernel (see figure 6). We also find that GFSVI requires a larger number of measurement points to capture the behavior of a rougher prior (see figure 7).
Figure 5: Our method (GFSVI) with a Matern-1/2 Gaussian process (GP) prior accurately approximates the exact GP posterior unlike the function-space prior baseline (FVI). Weight-space prior baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (MFVI, Laplace). The lower row is identical to the one in figure 2 in the main text and is reproduced here to make comparison easier.
@@
Figure 7: Our method (GFSVI) captures the smooth behavior of a Gaussian process (GP) prior with RBF covariance function even if the number of measurement points is small (M=10). However, in that setting GFSVI fails to reproduce the rough effect of a GP prior with a Matern-1/2 covariance function, and requires a larger amount of measurement points to do so (M=100).
@@
We find that GFSVI better captures the beliefs induced by the smooth RBF and rough Matern-1/2 Gaussian process priors compared to FVI (see Figures 9 and 10). Moreover, GFVSI both accurately fits the training data and shows greater uncertainty outside of its support relative to BNNs baselines with weight-space and function-space priors. Unlike for the toy data regression experiments where the GP posterior was the ground truth, the Laplace (approximate) GP posterior in Figures 9 and 10 only represents a possible approximation to the now in-tractable posterior (due to the softmax inverse link function). Thus the GP should not be considered as the ground truth nor as the optimal approximation in the classification setting, but is nevertheless useful to give a idea of the level of uncertainty a BNN with a GP prior should provide outside of the support of the data.
@@
Figure 10: Our method (GFSVI) with a Matern-1/2 Gaussian process (GP) prior accurately captures the rough decision boundary unlike the function-space baseline (FVI). Weight-space baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (TFSVI, Laplace) or underfit the data (MFVI).
@@
figure 11 compares GFSVI to the exact posterior across two different priors and three model architectures (details in A.3.1). We find that the BNN’s ability to incorporate the beliefs introduced by the GP prior depends on its size and activation function. When using piece-wise linear activations (ReLU), small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matern-1/2). By contrast, when using smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matern family, and Rational Quadratic in our experiments). We also analyzed how the number $M$ of measurement points affects performance. figures 7 and 14 show that capturing the properties of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
@@
A.4.2 Regression on tabular data
@@
A.4.3 Out-of-distribution detection with image data
We here show an additional plot from our out-of-distribution detection experiment with image data (details in A.3.5). figure 12 shows the (normalized) histograms of the entropy of the mean prediction produced by each model on the in-distribution (blue) and out-of-distribution (red) data sets considered in our OOD detection experiment. Methods which estimate the (regularized) KL-divergence in function-space (GFSVI, FVI and TFSVI) use the kmnist measurement distribution. We find that the entropy produced by GFSVI on in-distribution data highly peaks around 0 while the entropy produced from out-of-distribution data strongly concentrates around its maximum $\ln(10)$. GFSVI best partitions ID and OOD data based on predictive entropy improving over the function-space prior (FVI) and weight-space prior (TFSVI, MFVI, Laplace) BNN baselines (see table 2).
@@ A.4.3 Out-of-distribution detection with image data
A.4.4 Influence of measurement point distribution for image data.
We present additional results evaluating the influence of the measurement point distribution $\rho$ on the the performance of function-space inference methods when using high-dimensional image data. The measurement point distribution are described in section A.3.4. Just like in Rudner et al., 2022b , we find that the choice of measurement point distribution may highly influence the OOD detection accuracy. While the expected log-likelihood, accuracy and expected calibration error (ECE) of a model generally remains comparable across measurement point distributions, the OOD accuracy of GFSVI is greatly improved by using samples from kmnist to evaluate the (regularized) KL divergence. The measurement point distribution determines where the BNN is regularized and thus should be carefully selected especially for high dimensional data.
@@
A.4.5 Input distribution shift with rotated image data
We here provide an experiment evaluating our method’s (GFSVI) robustness in detecting input distribution shift. We expect the predictive uncertainty of a well-calibrated Bayesian model to be low for in-distribution data and to gradually increase as the input distribution shifts further away from the training data distribution. To test this property, we follow the setup by Sensoy et al., (2018); Rudner et al., 2022b and assume like the related work that increasing the rotation angle of images gradually increases the level of input "distribution shift". We report the mean and standard-deviation of the average mean predictive entropy of models fit on MNIST (LeCun et al.,, 2010) and Fashion MNIST (Xiao et al.,, 2017) for increasingly large angles of rotation of their respective test data partition. We find that GFSVI is confident (low predictive entropy) for images with small rotation angles, and that its predictive entropy increases with the angle. GFSVI therefore exhibits the expected behavior of a well-calibrated Bayesian model. We note that FVI, Laplace and MFVI tend to be under-confident (high predictive entropy) for small rotation angles, which might be a symptom of underfitting further supported by the results in table 2. Also, with the exception of TFSVI, the predictive entropy of baselines across different rotation angles is generally higher than the one produced by GFSVI.
@@ A.4.5 Input distribution shift with rotated image data
A.4.6 Additional plots for kernel eigenvalue decay
figure 14 shows a plot demonstrating the decay rate of the eigenvalues of RBF and Matern-1/2 kernels evaluated at points sampled uniformly over $\mathcal{X}$ The rate of decay of covariance operator’s eigenvalues gives important information about the smoothness of stationary kernels (Williams and Rasmussen,, 2006) and that increased smoothness of the kernel leads to faster decay of eigenvalues Santin and Schaback, (2016). For instance, RBF covariance operator eigenvalues decay at near exponential rate independent of the underlying measure (Belkin,, 2018) and Matern kernels eigenvalues decay polynomialy (Chen et al.,, 2021). We find that the kernel evaluated at points sampled from a uniform distribution over $\mathcal{X}$ share this same behavior (see figure 14).
@@ A.4.6 Additional plots for kernel eigenvalue decay
Figure 14: Mean eigenvalues of the Gram matrix obtained for different kernels and for varying length-scales over $10$ draws from a uniform distribution on $[-2,2]^{D}$. The mean eigenvalues are arranged in increasing order. The eigenvalues of the Gram matrix associated with the smooth RBF kernel decays much faster than those of the Matern-1/2. Furthermore, the eigenvalues decay at a slower rate in high dimensions (D=100).
A.4.7 Additional plots for choosing $\gamma$ in $D_{\text{KL}}^{\gamma}$
@@
Figure 16: The BNN’s covariance adaptation to the prior’s covariance rank depends on its activation function. BNNs fit with a RBF prior (full) show lower rank than with a Matern-1/2 (dotted).
Figure 17: $\gamma$ explicitly controls the magnitude of the regularized KL-divergence $D_{\text{KL}}^{\gamma}$. Rougher priors (Matern-1/2) require more measurement points to accurately estimate $D_{\text{KL}}^{\gamma}$ than smooth priors (RBF).

### v3

@@ C.8 Hardware
Appendix D Additional experimental results
@@
D.1 Qualitative uncertainty evaluation
@@
We further find that our method (GFSVI) provides strong regularization when the data generative process is noisy (see figure 4) and is more robust than FVI to situations where ones computational budget constrains the number of measurement points $M$ to be small (figure 8). In contrast to FVI, GFSVI accurately approximates the exact GP posterior under rough (Matérn-1/2) GP priors effectively incorporating prior knowledge defined by the GP prior to the inference process (see figure 2). Likewise, GFSVI adapts to the variability of the functions specified by the kernel (see figure 6). We also find that GFSVI requires a larger number of measurement points to capture the behavior of a rougher prior (see figure 7).
Figure 5: Our method (GFSVI) with an RBF Gaussian process (GP) prior accurately approximates the exact GP posterior unlike the function-space prior baseline (FVI). Weight-space prior baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (MFVI, Laplace). The lower row is identical to the one in figure 2 in the main text and is reproduced here to make comparison easier.
@@
Figure 7: Our method (GFSVI) captures the smooth behavior of a Gaussian process (GP) prior with RBF covariance function even if the number of measurement points is small (M=10). However, in that setting GFSVI fails to reproduce the rough effect of a GP prior with a Matérn-1/2 covariance function, and requires a larger amount of measurement points to do so (M=100).
@@
We find that GFSVI better captures the beliefs induced by the smooth RBF and rough Matérn-1/2 Gaussian process priors compared to FVI (see Figures 9 and 10). Moreover, GFVSI both accurately fits the training data and shows greater uncertainty outside of its support relative to BNNs baselines with weight-space and function-space priors. Unlike for the toy data regression experiments where the GP posterior was the ground truth, the Laplace (approximate) GP posterior in Figures 9 and 10 only represents a possible approximation to the now in-tractable posterior (due to the softmax inverse link function). Thus the GP should not be considered as the ground truth nor as the optimal approximation in the classification setting, but is nevertheless useful to give a idea of the level of uncertainty a BNN with a GP prior should provide outside of the support of the data.
@@
Figure 10: Our method (GFSVI) with a Matérn-1/2 Gaussian process (GP) prior accurately captures the rough decision boundary unlike the function-space baseline (FVI). Weight-space baselines do not provide a straight-forward mechanism to incorporate prior assumptions regarding the functions generated by BNNs and underestimate the epistemic uncertainty (TFSVI, Laplace) or underfit the data (MFVI).
@@
figure 11 compares GFSVI to the exact posterior across two different priors and three model architectures (details in C.1). We find that the BNN’s ability to incorporate the beliefs introduced by the GP prior depends on its size and activation function. When using piece-wise linear activations (ReLU), small models are prone to underfitting for smooth priors (RBF), and to collapsing uncertainty for rough priors (Matérn-1/2). By contrast, when using smooth activations (Tanh), smaller models suffice, and they are compatible with most standard GP priors (the results shown in figure 11 extend to RBF, Matérn family, and Rational Quadratic in our experiments). We also analyzed how the number $M$ of measurement points affects performance. figures 7 and 17 show that capturing the properties of rough GP priors and estimating $D_{\text{KL}}^{\gamma}$ with these priors requires larger $M$.
@@
D.2 Regression on tabular data
@@
D.3 Variational measure evaluation
table 7 evaluates our inference method by comparing samples drawn from the exact posterior (where computationally feasible) with the approximate posterior obtained with our method (GFSVI). We follow the setup by Wilson et al. [2022] and we compute the average per-sample Wasserstein-2 metric samples drawn from a GP posterior with RBF kernel evaluated at the test points, and samples from the approximate posterior of GFSVI, sparse GP and FVI evaluated at the same points and with the same prior. More details are provided in section C.6. We find that GFSVI approximates the exact posterior more accurately that FVI, obtaining a higher mean rank, but worse than the gold standard sparse GP, which demonstrates to be most accurate.
Dataset Boston Concrete Energy Wine Yacht Mean rank
D.4 Out-of-distribution detection with image data
We here show an additional plot from our out-of-distribution detection experiment with image data (details in C.5). figure 12 shows the (normalized) histograms of the entropy of the mean prediction produced by each model on the in-distribution (blue) and out-of-distribution (red) data sets considered in our OOD detection experiment. Methods which estimate the (regularized) KL-divergence in function-space (GFSVI, FVI and TFSVI) use the kmnist measurement distribution. We find that the entropy produced by GFSVI on in-distribution data highly peaks around 0 while the entropy produced from out-of-distribution data strongly concentrates around its maximum $\ln(10)$. GFSVI best partitions ID and OOD data based on predictive entropy improving over the function-space prior (FVI) and weight-space prior (TFSVI, MFVI, Laplace) BNN baselines (see table 2).
@@ D.4 Out-of-distribution detection with image data
D.5 Influence of measurement point distribution for image data
We present additional results evaluating the influence of the measurement point distribution $\rho$ on the the performance of function-space inference methods when using high-dimensional image data. The measurement point distribution are described in section C.4. Just like in Rudner et al. [2022b], we find that the choice of measurement point distribution may highly influence the OOD detection accuracy. While the expected log-likelihood, accuracy and expected calibration error (ECE) of a model generally remains comparable across measurement point distributions, the OOD accuracy of GFSVI is greatly improved by using samples from kmnist to evaluate the (regularized) KL divergence. The measurement point distribution determines where the BNN is regularized and thus should be carefully selected especially for high dimensional data.
@@ D.5 Influence of measurement point distribution for image data
D.6 Input distribution shift with rotated image data
We here provide an experiment evaluating our method’s (GFSVI) robustness in detecting input distribution shift. We expect the predictive uncertainty of a well-calibrated Bayesian model to be low for in-distribution data and to gradually increase as the input distribution shifts further away from the training data distribution. To test this property, we follow the setup by Sensoy et al. [2018], Rudner et al. [2022b] and assume like the related work that increasing the rotation angle of images gradually increases the level of input "distribution shift". We report the mean and standard-deviation of the average mean predictive entropy of models fit on MNIST [LeCun et al., 2010] and Fashion MNIST [Xiao et al., 2017] for increasingly large angles of rotation of their respective test data partition. We find that GFSVI is confident (low predictive entropy) for images with small rotation angles, and that its predictive entropy increases with the angle. GFSVI therefore exhibits the expected behavior of a well-calibrated Bayesian model. We note that FVI, Laplace and MFVI tend to be under-confident (high predictive entropy) for small rotation angles, which might be a symptom of underfitting further supported by the results in table 2. Also, with the exception of TFSVI, the predictive entropy of baselines across different rotation angles is generally higher than the one produced by GFSVI.
@@ D.6 Input distribution shift with rotated image data
D.7 Example of model misspecification with Wild et al. [2022b]
Figure 14: Example of model misspecification when using a periodic GP prior with baseline GWI [Wild et al., 2022b] that does not occur with our method (GFSVI). In GWI, only the posterior covariance is periodic, while the neural network parameterizing the posterior mean results in a function that does not capture the beliefs carried by the (periodic) prior. In contrast, our method accurately captures the GP prior’s beliefs and yields a (locally) periodic function.
figure 14 shows an example of model misspecification when using a periodic GP prior with the baseline GWI [Wild et al., 2022b]. As can be seen in the left panel of the figure, this problem does not occur with our method (GFSVI). While the posterior covariance in GWI reflects the periodicity of the prior, the neural network parametrizing the posterior mean does not result in a periodic function, i.e., the mean does not capture the beliefs specified by the periodic GP prior. In contrast, our method accurately captures the GP prior’s beliefs and yields a (locally) periodic function.
D.8 Convergence speed on UCI data
Table 9: Training time of our method GFSVI and baselines MFVI [Blundell et al., 2015] and TFSVI [Rudner et al., 2022b] on the boston UCI dataset.
GFSVI (ours) TFSVI MFVI
Time (s) 44.15 $\pm$ 1.56 36.36 $\pm$ 0.90 38.38 $\pm$ 10.80
table 9 shows the training time of our method and baselines MFVI [Blundell et al., 2015] and TFSVI [Rudner et al., 2022b] on the boston dataset using $M=100$ context points averaged over $5$ cross-validation splits, as well as figure 15 showing the convergence of the validation expected log-likelihood on the boston dataset. Our method converges in more steps than the TFSVI. GFSVI typically takes more time/steps to train that TFSVI as it additionally needs to adapt its features to the beliefs specified by the Gaussian process prior.
Figure 15: Validation expected log-likelihood of our method (GFSVI) and baselines TFSVI and MFVI. GFSVI (ours) converges on the boston dataset in slightly more steps than TFSVI but in fewer than MFVI.
Figure 16: Validation expected log-likelihood of our method (GFSVI) and baselines TFSVI and MFVI on MNIST. GFSVI (ours) converges in slightly more steps than TFSVI but in fewer than MFVI.
D.9 Additional plots for kernel eigenvalue decay
figure 17 shows a plot demonstrating the decay rate of the eigenvalues of RBF and Matérn-1/2 kernels evaluated at points sampled uniformly over $\mathcal{X}$ The rate of decay of covariance operator’s eigenvalues gives important information about the smoothness of stationary kernels [Williams and Rasmussen, 2006] and that increased smoothness of the kernel leads to faster decay of eigenvalues Santin and Schaback [2016]. For instance, RBF covariance operator eigenvalues decay at near exponential rate independent of the underlying measure [Belkin, 2018] and Matérn kernels eigenvalues decay polynomialy [Chen et al., 2022]. We find that the kernel evaluated at points sampled from a uniform distribution over $\mathcal{X}$ share this same behavior (see figure 17).
@@ D.9 Additional plots for kernel eigenvalue decay
Figure 17: Mean eigenvalues of the Gram matrix obtained for different kernels and for varying length-scales over $10$ draws from a uniform distribution on $[-2,2]^{D}$. The mean eigenvalues are arranged in increasing order. The eigenvalues of the Gram matrix associated with the smooth RBF kernel decays much faster than those of the Matérn-1/2. Furthermore, the eigenvalues decay at a slower rate in high dimensions (D=100).
D.10 Additional plots for choosing $\gamma$ in $D_{\text{KL}}^{\gamma}$
@@
Figure 19: The BNN’s covariance adaptation to the prior’s covariance rank depends on its activation function. BNNs fit with a RBF prior (full) show lower rank than with a Matérn-1/2 (dotted).
Figure 20: $\gamma$ explicitly controls the magnitude of the regularized KL-divergence $D_{\text{KL}}^{\gamma}$. Rougher priors (Matérn-1/2) require more measurement points to accurately estimate $D_{\text{KL}}^{\gamma}$ than smooth priors (RBF).